# Covariance of Scattering Amplitudes from Counting Carefully

### v2

@@ 1 Introduction
The invariance of physical amplitudes under certain transformations known as field redefinitions is a classical result in quantum field theory. When we compute an observable within the path integral formalism, such as a scattering amplitude, the ‘field’ $\phi(x)$ may be treated as an integration variable. The invariance of observables under field redefinitions is no different than the invariance of an integral under a change of variables Arzt1995:Covariance ; CHISHOLM1961:Fieldredefinitions ; KAMEFUCHI1961:FieldRedefinitions .
Scattering amplitudes are obtained from $\mathcal{S}$-matrix elements using the Lehmann–Symanzik –Zimmermann (LSZ) formula  LSZ:1955
@@
To resolve this well known ambiguity, it is valuable to develop computational techniques that are manifestly covariant, and to try to understand the structure of Lagrangians in terms of covariant objects. In the context of physics Beyond the Standard Model (BSM), the use of effective field theories (EFTs) makes the issue particularly pertinent, as the BSM effects are captured via either of the Standard Model Effective Field Theory (SMEFT) or the Higgs Effective Field Theory (HEFT), which differ from one another in how the scalar sector is parametrized. A SMEFT Lagrangian may always be recast into a HEFT Lagrangian, but the inverse map is singular, meaning there exist BSM scenarios that can match only to HEFT, not SMEFT Cohen:2020xca ; Falkowski2019:WhichEFT ; Alonso_2016:GeometryScalar ; Alonso_2015:GeometryHiggs ; Helset_2020:GeoSmeft . This presents a need to study the physical amplitudes or equivalently the connected amputated functions $\mathcal{A}_{a_{1}\dots a_{n}}$ for arbitrary values of $n$ to identify whether a HEFT theory encodes the same physics as a SMEFT theory.
While covariance is a fundamental property of $\mathcal{A}_{a_{1}\dots a_{n}}$, it can be explicitly highlighted through the use of different formalisms, with the use of geometry to represent the Lagrangian emerging as a favored approach for distinguishing HEFT theories from SMEFT theories in recent years; by identifying the fields with maps into a manifold, the kinetic (two derivative) term naturally finds a description in terms of a (pseudo-)Riemannian metric $g_{ij}(\phi)\partial_{\mu}\phi^{i}\partial^{\mu}\phi^{j}$, which in turn allows for its contributions to the amputated connected function to be expressed in terms of covariant geometric objects  Alonso_2015:GeometryHiggs ; Alonso_2016:GeometryScalar ; Alonso_2016:SigmaModels ; Cohen:2020xca . The Levi-Civita connection defined by the metric tensor makes the derivation of covariant Feynman rules in geometric theories particularly simple Apostolos2020:Covariance ; alminawi2025scalaramps .
@@
In more detail, in the diagrammatic approach the amputated connected functions are obtained via Wick contractions of 1PI vertices with propagators. The 1PI functions themselves are not individually covariant in general. Consequently, in the diagrammatic approach the covariance of connected functions is obscured behind a series of cancellations that must occur between the different Feynman diagrams. This structure has been examined across different contexts, using tools from combinatorics in  FIGUEROA:HopfAlgebra ; Kreimer:Combinatorics2016 ; Balduf:Combinatorics2020 . Restoring covariance at the level of 1PI functions was demonstrated using the path integral formalism in  Apostolos2020:Covariance ; HONERKAMP1972:Multiloops , with the validity being showcased through computations in specific examples relevant to pion physics in  HONERKAMP1974:OnShell ; ECKER1972:CovariantPerturbation .
Computing $\mathcal{A}_{a_{1}\dots a_{n}}$ for high values of $n$, which is a challenging task, has been tackled by several previous works, which this paper will draw upon. Combinatorial approaches have provided an iterative result for the ”tree sums”  Kreimer:Combinatorics2016 ; Balduf:Combinatorics2020 . However, the result does not distinguish the different possible index contractions among identical vertices. The path integral formalism has been used to derive a recursive method for the computation of Feynman rules Cohen:2023Covariance ; Cohen:2024Geometry ; cohen2025geometricbuildingblocks , which produce the amputated functions $\mathcal{A}_{a_{1}\dots a_{n}}$ up to ”evanescent” terms, i.e terms that vanish on-shell. The formalism has also been used to derive manifestly covariant Feynman rules Apostolos2020:Covariance ; HONERKAMP1972:Multiloops .
@@
  • The existence of covariant Feynman rules and the validity of using them to compute amputated connected functions is shown for any $n$ at the tree level, the efficiency of this approach is demonstrated in alminawi2025scalaramps using covariant Feynman rules up to $n=10$. The validity of the results for an arbitrary number of derivatives also makes them a valuable asset for the application of geometric techniques to higher derivatives through the use of jet bundles Alminawi_2024:JetBundles or functional geometry  Apostolos2020:Covariance ; cohen2025geometricbuildingblocks ; Helset2022:Geometry ; Cohen:2024Geometry ; Cheung2022:GeometryKinematics ; Cohen:2023Covariance
@@ 2 Covariance as a combinatoric property
The covariance of the connected amputated $n$-point function on-shell, particularly at the tree level, can easily be showcased through the use of generating functionals as done in Cohen:2023Covariance ; Cohen2023:Letter . However, in practice the derivation of $\mathcal{A}_{a_{1}\dots a_{n}}$ often follows a different path, starting from the effective action $\Gamma[\phi]$, we can identify 1PI functions as the functional derivatives of the effective action
@@
For an arbitrary $n$ the transformation behavior of the 1PI function is encoded by the Faà di Bruno formula Comtet_Combinatorics ; Riordan_Combinatorial
@@
where $B_{n,k}$ are the partial exponential Bell polynomials defined by
@@
Connected $n$-point functions $\mathcal{A}_{a_{1}\dots a_{n}}$ can be constructed by connecting 1PI functions with $n\geq 3$ to one another using propagators $D_{F}$ as internal lines, which will be referred to as ”gluing” in this paper, and forming all possible tree level topologies as shown in Fig.  1
@@
To identify the necessary cancellations between the different diagrams, it is crucial to count the number of topologies correctly. At the tree level, the number of channels is equal to $n!/|\text{Aut}|$ where $n$ is the number of external legs and $|\text{Aut}|$ is the order of the automorphism group of the diagram, which accounts for the symmetries of a given diagram, for example consider the first two diagrams in Fig  2, in the first diagram swapping any of the legs with each other results in the same diagram, thus $|\text{Aut}|=4!$ and the number of diagrams $n!/|\text{Aut}|=1$. For the second diagram notice that swapping $a_{1}\leftrightarrow a_{2}$, $a_{3}\leftrightarrow a_{4}$ or $\{a_{1},a_{2},b_{1}\}\leftrightarrow\{a_{3},a_{4},b_{2}\}$ all leave the diagram unchanged, each swap has order two, thus $|\text{Aut}|=8$ and thus the number of distinct diagrams is $n!/|\text{Aut}|=3$, with all three diagrams shown in the figure. Identifying the full set of Feynman diagrams and their symmetries for a given number of external points $n$ is a combinatorial and graph theory problem, whose solution is not known to the best of my knowledge, though several similar sequences are well known and thoroughly studied as mentioned in the following section.
@@ 3 Counting Tree Graphs
The construction of the connected function from 1PI vertices requires the identification of all the possible tree diagrams as well as the number of topologies associated with each diagram. For $n\geq 4$ external legs there is more than one type of diagram contributing to the connected function
@@
Figure 2: Diagrams contributing to four point connected function
@@ 3.3 Counting topologies
The degree of a node refers to the number of edges to which it is connected. A node of degree 1 is called a leaf, which corresponds to an external leg in a Feynman diagram. Trees in which all nodes have degree 1 or degree $\geq 3$ are special, they are referred to as series reduced trees (or phylogenetic trees when rooted [^2] A root is simply a specification of an external node as a starting point in a tree graph) ElveyPrice2020:PhylogeneticTrees .
The set of tree level Feynman diagrams is precisely that of series reduced trees, which can be proven by induction; starting with four nodes for the base case, it is clear that there is only one tree graph, which is the one shown in the second line of Fig. 5 and similarly there is only one Feynman diagram. Feynman diagrams are unrooted trees, a Feynamn diagram with $n$ external legs and $k$ vertices corresponds to a phylogenetic tree of type (n2,k)(n-2,k) [^3] The convention for Ward numbers states that an (n2,k)(n-2,k) phylogenetic tree is a rooted tree with n1n-1 external nodes and $k$ internal nodes, the Ward numbers Ward1934:PhylogeneticTrees counting such trees are generated by the recursion relation
@@
The total number of rooted series reduced diagrams corresponds to the sum over Ward numbers for a fixed number of external nodes $n$, it can be found in the Online Enyclopedia of Integer Sequences as OEIS A000311 oeis , which reads
@@
The Ward numbers defined by the recursion relation are given by OEIS A269939 oeis which split the number of rooted trees with $n+1$ external nodes based on the number of internal nodes $k$
@@
As mentioned earlier there are multiple combinations of vertices at the same level of gluing that contribute to the connected function such as in Fig. 3, which creates a need for going beyond the Ward numbers. A natural refinement can be obtained by examining a multivariate polynomial. The multivariate Ward polynomial is obtained by labeling each vertex by $a_{i-2}$ where $i$ is the number of legs HAIMAN1989:LagrangeInversion . For the 6 point case, corresponding to $W(4,k)$ it gives
@@
The multivariate Ward polynomial bears close resemblance to the Bell polynomial  Comtet_Combinatorics ; Riordan_Combinatorial . In fact upon close examination it is clear that
@@
The multivariate polynomial is known, its coefficients (up to signs) can be found as OEIS Sequence A134685  oeis , the first few polynomials are given by the table below
@@ 4 $a_{4}$ $10a_{2}^{2}+15a_{3}a_{1}$ $105a_{2}a_{1}^{2}$ $105a_{1}^{4}$
To recover the Ward polynomial simply evaluate $W_{n,k}(1,x,x^{2},\dots)$ and it is trivial to see that the right hand side fulfills the recursion relation defining the Ward numbers. The full tree sums given by $W_{n}$ can be obtained by evaluating $W_{n,k}(1,1,\dots)$. The multivariate Ward polynomial describes the Faà di Bruno algebra FIGUEROA:HopfAlgebra ; Connes2000:Hopf1 and can be related to Feynman graphs  Connes2001:Renormalization ; Abdesselam2002:Feynman .
@@
as found in Bern_2008 , which corresponds to the case $W(n,n)$. If the number of edges is denoted by $m$ and the number of diagrams by $f_{n}$, then the sequence is given by

### v3

@@ 1 Introduction
The invariance of physical amplitudes under certain transformations known as field redefinitions is a classical result in quantum field theory. When we compute an observable within the path integral formalism, such as a scattering amplitude, the ‘field’ $\phi(x)$ may be treated as an integration variable. The invariance of observables under field redefinitions is no different than the invariance of an integral under a change of variables Arzt (1995); Chisholm (1961); Kamefuchi et al. (1961).
Scattering amplitudes are obtained from $\mathcal{S}$-matrix elements using the Lehmann–Symanzik –Zimmermann (LSZ) formula  Lehmann et al. (1955)
@@
To resolve this well known ambiguity, it is valuable to develop computational techniques that are manifestly covariant, and to try to understand the structure of Lagrangians in terms of covariant objects. In the context of physics Beyond the Standard Model (BSM), the use of effective field theories (EFTs) makes the issue particularly pertinent, as the BSM effects are captured via either of the Standard Model Effective Field Theory (SMEFT) or the Higgs Effective Field Theory (HEFT), which differ from one another in how the scalar sector is parametrized. A SMEFT Lagrangian may always be recast into a HEFT Lagrangian, but the inverse map is singular, meaning there exist BSM scenarios that can match only to HEFT, not SMEFT Cohen et al. (2021); Falkowski and Rattazzi (2019); Alonso et al. (2016b, a); Helset et al. (2020). This presents a need to study the physical amplitudes or equivalently the connected amputated functions $\mathcal{A}_{a_{1}\dots a_{n}}$ for arbitrary values of $n$ to identify whether a HEFT theory encodes the same physics as a SMEFT theory.
While covariance is a fundamental property of $\mathcal{A}_{a_{1}\dots a_{n}}$, it can be explicitly highlighted through the use of different formalisms, with the use of geometry to represent the Lagrangian emerging as a favored approach for distinguishing HEFT theories from SMEFT theories in recent years; by identifying the fields with maps into a manifold, the kinetic (two derivative) term naturally finds a description in terms of a (pseudo-)Riemannian metric $g_{ij}(\phi)\partial_{\mu}\phi^{i}\partial^{\mu}\phi^{j}$, which in turn allows for its contributions to the amputated connected function to be expressed in terms of covariant geometric objects  Alonso et al. (2016a, b, c); Cohen et al. (2021). The Levi-Civita connection defined by the metric tensor makes the derivation of covariant Feynman rules in geometric theories particularly simple Finn et al. (2020); Alminawi et al. (2025).
@@
In more detail, in the diagrammatic approach the amputated connected functions are obtained via Wick contractions of 1PI vertices with propagators. The 1PI functions themselves are not individually covariant in general. Consequently, in the diagrammatic approach the covariance of connected functions is obscured behind a series of cancellations that must occur between the different Feynman diagrams. This structure has been examined across different contexts, using tools from combinatorics in  FIGUEROA and GRACIA-BONDÍA (2005); Kreimer and Yeats (2017); Balduf (2020). Restoring covariance at the level of 1PI functions was demonstrated using the path integral formalism in  Finn et al. (2020); Honerkamp (1972), with the validity being showcased through computations in specific examples relevant to pion physics in  Honerkamp et al. (1974); Ecker and Honerkamp (1972).
Computing $\mathcal{A}_{a_{1}\dots a_{n}}$ for high values of $n$, which is a challenging task, has been tackled by several previous works, which this paper will draw upon. Combinatorial approaches have provided an iterative result for the ”tree sums”  Kreimer and Yeats (2017); Balduf (2020). However, the result does not distinguish the different possible index contractions among identical vertices. The path integral formalism has been used to derive a recursive method for the computation of Feynman rules Cohen et al. (2024, 2025b, 2025a), which produce the amputated functions $\mathcal{A}_{a_{1}\dots a_{n}}$ up to ”evanescent” terms, i.e terms that vanish on-shell. The formalism has also been used to derive manifestly covariant Feynman rules Finn et al. (2020); Honerkamp (1972).
@@
  • The existence of covariant Feynman rules and the validity of using them to compute amputated connected functions is shown for any $n$ at the tree level, the efficiency of this approach is demonstrated in Alminawi et al. (2025) using covariant Feynman rules up to $n=10$. The validity of the results for an arbitrary number of derivatives also makes them a valuable asset for the application of geometric techniques to higher derivatives through the use of jet bundles Alminawi et al. (2024) or functional geometry  Finn et al. (2020); Cohen et al. (2025a); Helset et al. (2022); Cohen et al. (2025b); Cheung et al. (2022); Cohen et al. (2024)
@@ 2 Covariance as a combinatoric property
The covariance of the connected amputated $n$-point function on-shell, particularly at the tree level, can easily be showcased through the use of generating functionals as done in Cohen et al. (2024, 2023). However, in practice the derivation of $\mathcal{A}_{a_{1}\dots a_{n}}$ often follows a different path, starting from the effective action $\Gamma[\phi]$, we can identify 1PI functions as the functional derivatives of the effective action
@@
For an arbitrary $n$ the transformation behavior of the 1PI function is encoded by the Faà di Bruno formula Comtet (1974); Riordan (1968)
@@
where $\psi^{\prime}=\frac{\delta\psi}{\delta\phi}$ and $B_{n,k}$ are the partial exponential Bell polynomials defined by
@@
Under an arbitrary field redefinition $\phi\to\phi+f(\phi)$ the new Feynman rules emerging will be dependent on terms from Feynman rules with fewer legs, for example the four point Feynman rule with suppressed indices after the transformation is given by
$\frac{\delta^{4}}{\delta\phi^{4}}\Gamma[\psi(\phi)]=\frac{\delta^{4}\Gamma[\psi]}{\delta\psi^{4}}\psi^{\prime 4}+6\frac{\delta^{3}\Gamma[\psi]}{\delta\psi^{3}}\psi^{\prime 2}\psi^{\prime\prime}+3\frac{\delta^{2}\Gamma[\psi]}{\delta\psi^{2}}\psi^{\prime\prime 2}+4\frac{\delta^{2}\Gamma[\psi]}{\delta\psi^{2}}\psi^{\prime}\psi^{\prime\prime\prime}$ (8)
after imposing the vacuum condition.
There are two terms proportional to the inverse propagator, namely $3\frac{\delta^{2}\Gamma[\psi]}{\delta\psi^{2}}\psi^{\prime\prime 2}$ and $4\frac{\delta^{2}\Gamma[\psi]}{\delta\psi^{2}}\psi^{\prime}\psi^{\prime\prime\prime}$, the on-shell condition is only relevant for the second term, to understand this consider the definitions of $\phi^{2}$ and $\phi^{3}$
$\phi^{2}(x)=\int\frac{d^{d}p_{1}}{(2\pi)^{d}}\frac{d^{d}p_{2}}{(2\pi)^{d}}\phi(p_{1})\phi(p_{2})e^{i(p_{1}+p_{2})x}$ (9)
and
$\phi^{3}(x)=\int\frac{d^{d}p_{1}}{(2\pi)^{d}}\frac{d^{d}p_{2}}{(2\pi)^{d}}\frac{d^{d}p_{3}}{(2\pi)^{d}}\phi(p_{1})\phi(p_{2})\phi(p_{3})e^{-i(p_{1}+p_{2}+p_{3})x}$ (10)
the term $3\frac{\delta^{2}\Gamma[\psi]}{\delta\psi^{2}}\psi^{\prime\prime 2}$ picks up factors from $\phi^{2}$ as indicated by $\psi^{\prime\prime}$ and the coefficient 3, which indicates the three independent momenta combinations $p_{1}+p_{2},p_{1}+p_{3}$ and $p_{1}+p_{4}$, thus the term is of the form $\sim(p_{i}+p_{j})^{2}-m^{2}$ which does not vanish on-shell.
On the other hand if we consider the term $4\frac{\delta^{2}\Gamma[\psi]}{\delta\psi^{2}}\psi^{\prime}\psi^{\prime\prime\prime}$ it either picks up momenta from $\phi$ as indicated by $\psi^{\prime}$ or from $\phi^{3}$ through $\psi^{\prime\prime\prime}$, with the coefficient 4 indicating that there are 4 independent choices $p_{1},p_{2},p_{3}$ and $p_{4}$, using the fact that $p_{1}+p_{2}+p_{3}=-p_{4}$ by momentum conservation, it is clear that $\phi^{3}$ contributes in the same way as $\phi$ in the four point Feynman rule, thus the contribution has the form $\sim p_{i}^{2}-m^{2}$ which vanishes on-shell with $p_{i}^{2}=m^{2}$
The connected function is referred to as covariant if under the redefinition $\phi^{a_{i}}\to\psi^{b_{i}}(\phi)$ it transforms as
$\mathcal{A}_{a_{1}\dots a_{n}}\to\frac{\delta\psi^{b_{1}}}{\delta\phi^{a_{1}}}\frac{\delta\psi^{b_{2}}}{\delta\phi^{a_{2}}}\dots\frac{\delta\psi^{b_{n}}}{\delta\phi^{a_{n}}}\mathcal{A}_{b_{1}\dots b_{n}}$ (11)
with suppressed indices this corresponds to all terms proportional to $\psi^{\prime\prime},\psi^{\prime\prime\prime}\dots\psi^{(n)}$ vanishing.
Connected $n$-point functions $\mathcal{A}_{a_{1}\dots a_{n}}$ can be constructed by connecting 1PI functions with $n\geq 3$ to one another using propagators $D_{F}$ as internal lines, which will be referred to as ”gluing” in this paper, and forming all possible tree level topologies as shown in Fig. 1
@@
To identify the necessary cancellations between the different diagrams, it is crucial to count the number of topologies correctly. At the tree level, the number of channels is equal to $n!/|\text{Aut}|$ where $n$ is the number of external legs and $|\text{Aut}|$ is the order of the automorphism group of the diagram, which accounts for the symmetries of a given diagram, for example consider the first two diagrams in Fig. 2, in the first diagram swapping any of the legs with each other results in the same diagram, thus $|\text{Aut}|=4!$ and the number of diagrams $n!/|\text{Aut}|=1$. For the second diagram notice that swapping $a_{1}\leftrightarrow a_{2}$, $a_{3}\leftrightarrow a_{4}$ or $\{a_{1},a_{2},b_{1}\}\leftrightarrow\{a_{3},a_{4},b_{2}\}$ all leave the diagram unchanged, each swap has order two, thus $|\text{Aut}|=8$ and thus the number of distinct diagrams is $n!/|\text{Aut}|=3$, with all three diagrams shown in the figure. Identifying the full set of Feynman diagrams and their symmetries for a given number of external points $n$ is a combinatorial and graph theory problem, whose solution is not known to the best of my knowledge, though several similar sequences are well known and thoroughly studied as mentioned in the following section.
@@ 3 Counting Tree Graphs
The construction of the connected function from 1PI vertices requires the identification of all the possible tree diagrams as well as the number of topologies associated with each diagram. For $n\geq 4$ external legs there is more than one type of diagram contributing to the connected function, assigning arbitrary labels to the particles, using $a_{i}$ for external legs and $b_{i}$ for internal legs one finds
@@
Figure 2: Diagrams contributing to four point connected function. Labels correspond to the type of particle at each leg and their respective momenta to indicate the different channels
@@ 3.3 Counting topologies
The degree of a node refers to the number of edges to which it is connected. A node of degree 1 is called a leaf or an external node, which corresponds to an external leg in a Feynman diagram. Trees in which all nodes have degree 1 or degree $\geq 3$ are special, they are referred to as series reduced trees or as phylogenetic trees when considering rooted [^2] A root is simply a specification of an external node as a starting point in a tree graph trees.Elvey Price and Sokal (2020).
The set of tree level Feynman diagrams is precisely that of series reduced trees, which can be proven by induction; starting with four nodes for the base case, it is clear that there is only one tree graph, which is the one shown in the second line of Fig. 5 and similarly there is only one Feynman diagram. Feynman diagrams are unrooted trees, a Feynamn diagram with $n$ external legs and $k$ vertices corresponds to a phylogenetic tree of type (n2,k)(n-2,k) [^3] The convention for Ward numbers states that an (n2,k)(n-2,k) phylogenetic tree is a rooted tree with n1n-1 external nodes and $k$ internal nodes, the Ward numbers Ward (1934) counting such trees are generated by the recursion relation
@@
The total number of rooted series reduced diagrams corresponds to the sum over Ward numbers for a fixed number of external nodes $n$, it can be found in the Online Enyclopedia of Integer Sequences as OEIS A000311 OEIS Foundation Inc. , which reads
@@
The Ward numbers defined by the recursion relation are given by OEIS A269939  [^4] The table in the reference starts at n=1 corresponding to a graph with 2 external nodes OEIS Foundation Inc. , adapting the table to the case of Feynman diagrams yields
@@
As mentioned earlier there are multiple combinations of vertices at the same level of gluing that contribute to the connected function such as in Fig. 3, which creates a need for going beyond the Ward numbers. A natural refinement can be obtained by examining a multivariate polynomial. The multivariate Ward polynomial is obtained by labeling each vertex by $a_{i-2}$ where $i$ is the number of legs Haiman and Schmitt (1989). For the 6 point case, corresponding to $W(4,k)$ it gives
@@
The multivariate Ward polynomial bears close resemblance to the Bell polynomial  Comtet (1974); Riordan (1968). In fact upon close examination it is clear that
@@
The multivariate polynomial is known, its coefficients (up to signs) can be found as OEIS Sequence A134685  OEIS Foundation Inc. , the first few polynomials are given by the table below
@@ 4 $a_{4}$ $10a_{2}^{2}+15a_{3}a_{1}$ $105a_{2}a_{1}^{2}$ $105a_{1}^{4}$
to recover the Ward polynomial simply evaluate $W_{n,k}(1,x,x^{2},\dots)$ and it is trivial to see that the right hand side fulfills the recursion relation defining the Ward numbers. The full tree sums given by $W_{n}$ can be obtained by evaluating $W_{n,k}(1,1,\dots)$. The multivariate Ward polynomial describes the Faà di Bruno algebra FIGUEROA and GRACIA-BONDÍA (2005); Connes and Kreimer (2000) and can be related to Feynman graphs  Connes and Kreimer (2001); Abdesselam (2002).
@@
as found in Bern et al. (2008), which corresponds to the case $W(n,n)$. If the number of edges is denoted by $m$ and the number of diagrams by $f_{n}$, then the sequence is given by